# MultiMAE-DER: Multimodal Masked Autoencoder for Dynamic Emotion Recognition

### v1

@@ I Introduction
With the development of deep learning technology, convolutional neural networks (CNN) have shown excellent results in static emotion recognition, especially in facial expression recognition (FER) tasks, as seen in models like RTCNN [1], DeepEmotion [2], and PAtt-Lite [3]. Dynamic emotion recognition (DER) is one of the important fields of affective computing. Psychologists and Neuroscientists have been deeply exploring this field for a long time, while constructing a series of quantitative rules for recognizing emotional characteristics.
@@
In the field of computer vision, the introduction of Masked Autoencoder (MAE) strongly indicated that self-supervised learning would shine in the field of vision such as ImageMAE [10], VideoMAE [11], and AudioMAE [12]. On the other hand, while single-modal (visual or audio) inputs have shown good results in dynamic emotion recognition, as shown in models like HuBERT [13] and MAE-DFER [14], their performance has not yet reached the effect of real human emotion recognition. Therefore, multimodal (Visual-Audio) input has become a desirable development trend, as demonstrated by models like AV-LSTM [15] and MSAF [16].
@@ II-A Dynamic Emotion Recognition
Dynamic emotion recognition is a significant challenge in the field of affective computing that requires the utilization of multimodal feature extraction, context semantic analysis, and causal reasoning techniques. In recent years, researchers have addressed the feature extraction task from multimodal using transfer learning by combining pre-trained models for video facial expression feature extraction and audio tone emotion feature extraction, such as MSAF [16] and CFN-SR [17]. However, this model architecture fails to establish cross-domain associative features between visual and audio data, leading to the loss of highly correlated spatio-temporal feature dimensions in the emotion representation information. Therefore, we propose a MultiMAE-DER framework, a unified feature extraction pre-trained model that will be employed for visual and audio multimodal data to find the associated features in visual-audio spatio-temporal emotion representation information.
@@ II-B Masked Autoencoder
Masked Autoencoder is a variant of denoising autoencoder that consists of a pair of encoder and decoder. In ImageMAE [10], that is designed as an asymmetric structure to reconstruct normalized pixels. VideoMAE [11] follows the design approach of ImageMAE [10] but extends the 2D image input to 3D video input and incorporates a joint spatio-temporal self-attention mechanism to replace the vanilla Vision Transformer [18] spatial self-attention mechanism.
For downstream tasks, the decoder of the masked autoencoder is discarded, and only the pre-trained encoder is used for supervised learning and fine-tuning on the downstream task dataset. As a standout in self-supervised learning, our proposal will continue to inherit this characteristic and utilize the pre-trained encoder model to fine-tune multimodal dynamic emotion data, aiming for improved emotion recognition performance.
@@ IV-C Results
TABLE I: Results on RAVDESS. SSL: self-supervised learning method or not. UAR: unweighted average recall. WAR: weighted average recall. Underlined: the best supervised result. Bold: the best result.
@@
TABLE II: Results on CREMA-D. SSL: self-supervised learning method or not. UAR: unweighted average recall. WAR: weighted average recall. Underlined: the best supervised result. Bold: the best result.
@@
TABLE III: Results on IEMOCAP. SSL: self-supervised learning method or not. UAR: unweighted average recall. WAR: weighted average recall. Bold: the best result.
@@
Upon analyzing the outcomes across the three datasets, we observed that MultiMAE-DER exhibits negligible discrepancies in emotion recognition results between strategy 3 (FFLS) and strategy 4 (FSLF). This aligns with expectations, given that their distinctions lie solely in the sorting of multimodal sequences. Furthermore, the outcomes of these two strategies surpass those of other strategies, demonstrating enhancements ranging from 2% to 8% in WAR on the RAVDESS [19], 1% to 5% on the CREMA-D [20], and 1% to 3% on IEMOCAP [21].
In addition, in the case of single-modal input, MultiMAE-DER can find that the results of only audio signals are often better than those of only video signals. This shows that the audio signal contains more distinct and independent dynamic representation information. At the same time, better results will be produced compared to multi-modal input. It can be seen that the MultiMAE-DER model has the superiority in processing unified audio and video signals.
@@
V CONCLUSIONS AND FUTURE WORKS

### v2

@@ I Introduction
With the development of deep learning technology, convolutional neural networks (CNN) have shown excellent results in static emotion recognition, especially in facial expression recognition (FER) tasks, as seen in models like RTCNN [1], DeepEmotion [2], and PAtt-Lite [3]. Dynamic emotion recognition (DER) is one of the important fields of affective computing. Psychologists and neuroscientists have been exploring this field for a long time, while constructing a series of quantitative rules for recognizing emotional characteristics.
@@
In the field of computer vision, the introduction of Masked Autoencoder (MAE) strongly indicated that self-supervised learning would shine in the field of vision such as ImageMAE [10], VideoMAE [11], and AudioMAE [12]. On the other hand, while single-modal (visual or audio) inputs have shown good results in dynamic emotion recognition, as shown in models like HuBERT [13] and MAE-DFER [14], their performance has not yet reached the level of real human emotion recognition. Therefore, multimodal (Visual-Audio) input has become a desirable development trend, as demonstrated by models like AV-LSTM [15] and MSAF [16].
@@ II-A Dynamic Emotion Recognition
Dynamic emotion recognition is a significant challenge in the field of affective computing that requires the utilization of multimodal feature extraction, context semantic analysis, and causal reasoning techniques. In recent years, researchers have addressed the feature extraction task from multimodal using transfer learning by combining pre-trained models for video facial expression feature extraction and audio tone emotion feature extraction, such as MSAF [16] and CFN-SR [17]. However, the above-mentioned model architecture fails to establish cross-domain associative features between visual and audio data, leading to the loss of highly correlated spatio-temporal feature dimensions in the emotion representation information. Therefore, we propose a MultiMAE-DER framework, a unified feature extraction pre-trained model that will be employed for visual and audio multimodal data to find the associated features in visual-audio spatio-temporal emotion representation information.
@@ II-B Masked Autoencoder
Masked Autoencoder is a variant of denoising autoencoder that consists of an encoder and a decoder. In ImageMAE [10], that is designed as an asymmetric structure to reconstruct normalized pixels. VideoMAE [11] follows the design approach of ImageMAE [10] but extends the 2D image input to 3D video input and incorporates a joint spatio-temporal self-attention mechanism to replace the vanilla Vision Transformer [18] spatial self-attention mechanism.
For downstream tasks, the decoder of the masked autoencoder is discarded, and only the pre-trained encoder is used for supervised learning and fine-tuning on the downstream task dataset. As a leader in self-supervised learning, our proposal will continue to inherit this characteristic and utilize the pre-trained encoder model to fine-tune multimodal dynamic emotion data, aiming for improved emotion recognition performance.
@@ IV-C Results
TABLE I: Results on RAVDESS (7-class). SSL: self-supervised learning method or not. UAR: unweighted average recall. WAR: weighted average recall. Underlined: the best supervised result. Bold: the best result.
@@
TABLE II: Results on CREMA-D (6-class). SSL: self-supervised learning method or not. UAR: unweighted average recall. WAR: weighted average recall. Underlined: the best supervised result. Bold: the best result.
@@
TABLE III: Results on IEMOCAP (4-class). SSL: self-supervised learning method or not. UAR: unweighted average recall. WAR: weighted average recall. Bold: the best result.
@@
In addition, compared with the MultiMAE-DER single-modal model, the MultiMAE-DER multi-modal optimal strategy model (FSLF) is far superior in performance to the visual-only model and the audio-only model. The results show that the WAR on RAVDESS [19] is enhanced by 9.48% (83.61% vs. 74.13%) and 3.06% (83.61% vs. 80.55%) respectively, the WAR on CREMA-D [20] is enhanced by 1.53% (79.36% vs. 77.83%) and 0.91% (79.36% vs. 78.45%) respectively, and the WAR on IEMOCAP [21] is enhanced by 7.60% (63.73% vs. 56.13%) and 5.04% (63.73% vs. 58.69%) respectively.
Upon analyzing the outcomes across the three datasets, we observed that MultiMAE-DER exhibits negligible discrepancies in emotion recognition results between strategy 3 (FFLS) and 4 (FSLF). This aligns with expectations, given that their distinctions lie solely in the sorting of multimodal sequences. Furthermore, the effects of these two strategies surpass those of other strategies, demonstrating enhancements ranging from 2% to 8% in WAR on the RAVDESS [19], 1% to 5% on the CREMA-D [20], and 1% to 3% on IEMOCAP [21].
@@
V CONCLUSIONS AND FUTURE WORK